# PLS-complete problems with lexicographic cost functions: Max-$k$-SAT and Abelian Permutation Orbit Minimization

### v1

@@ 1.7 Our results
The proof, given in Section 3, follows the approach of Krentel [10, 9], who reduces from FLIP, but deviates from it in one crucial detail. With some additional work, we can show that $3$-SAT/2-FLIP with lexicographic payoffs is PLS-complete.
@@
Our final result answers an open question from Scheder and Tantow [16]
@@
We prove the theorem in Section 6 The proof follows from a reduction very similar to Buchheim and Jünger [1]. The permutations constructed in the reduction are commuting involutions (permutations of order 2) and thus the generated subgroup of the symmetric group is isomorphic to the additive group of $\mathbb{F}_{2}^{n}$, the vector space over $\mathbb{F}_{2}$.
We will give a second reduction, following the proof of Theorem 2 in [16], where all permutations are powers of one single permutation $\pi$. In other words, we prove the following theorem:
@@ 2 Krentel’s Approach and why it does not directly work for lexicographic Max-$k$-SAT
Krentel proved this to be PLS-complete[9] with the test circuit technique by reducing from FLIP. We sketch this proof again and show why it does not work for lexicographic weights. We are given a circuit $C$ with $n$ inputs and $m$ outputs. For the reduction, we imagine $n+1$ copies of $C$, called $C^{(0)},C^{(1)},\dots,C^{(n)}$. We want $C^{(0)}$ to compute $C(x)$ for the current input $x$, and each other $C^{(i)}$ ($1\leq i\leq n$) to compute $C(x\oplus\mathbf{e}_{i})$. For each circuit $C^{(i)}$ ($0\leq i\leq n$) we create a vector of $n$ input variables $\mathbf{x}^{(i)}=(x^{(i)}_{1},\dots,x^{(i)}_{n})$. We wish for $\mathbf{x}^{(i)}=\mathbf{x}^{(0)}\oplus\mathbf{e}_{i}$. For each circuit $i$ and gate $j$ we create a gate variable $g_{j}^{(i)}$. The variables involved in a gate can be consistent or inconsistent with the gate, and we call the gate correct or incorrect accordingly. We wish for all gates to be correct, but have to tolerate them being temporarily incorrect.
@@ 3 Lexicographic 4-SAT/2-Flip: Proof of Theorem 1
We have a switch variable $f$ (the needle in the above figure) that indicates whether TOP is considered the active circuit (if $f=1$) or BOTTOM is (if $f=0$). Finally, and this differs from [9], we have a global output register $\vec{o}=(o_{1},o_{2},\dots,o_{m})$.
@@
A difficulty. Simulating a correct evaluation of all gates is relatively easy to do with $k$-SAT/1-FLIP. The problem arises when all circuits are correct, but the payoff output of the inactive circuit is better than that of the active output. We need a mechanism that allows us to make the inactive circuit active (and vice versa) while improving the payoff (the total value of satisfied clauses). In [9] this is achieved by introducing clauses $(f\rightarrow v_{i}^{T})$ and $(\neg f\rightarrow v_{i}^{B})$, both of weight $2^{c-j}$ for some constant $c$. Making the BOTTOM active and TOP inactive (for example) will increase the total satisfied weight if and only if $\vec{v}^{B}>_{\rm lex}\vec{v}^{T}$. However, this only works if $(f\rightarrow v_{i}^{T})$ and $(\neg f\rightarrow v_{i}^{B})$ really have the same weight. If we insist on lexicographic weights and the first bit of both outputs is 0, say, and BOTTOM is active, then among $(f\rightarrow v_{1}^{T})$ and $(\neg f\rightarrow v_{1}^{B})$ only the first clause is satisfied, and it would never pay to make TOP active. This is where we need to introduce the global output and allow double flips.
@@ 6 Reduction to Abelian Polm—Proof of Theorem 4
We adapt a proof technique from [1] to reduce $4$-SAT/1-FLIP to Abelian Polm. Let $F$ be our $4$-CNF formula, let $x_{1},\dots,x_{n}$ be its variables, and let $C_{1},\dots,C_{m}$ be its clauses.
@@ 7 Reduction to Cyclic Group Polm—Proof of Theorem 5
Conversely, if $v$ is a local minimum, then $v$ has no $1$ at any bad-color position and thus $v$ encodes, as described above, a truth assignment $\alpha$ to the variables $x_{1},\dots,x_{n}$. Suppose $\alpha$ is not locally optimal for $F$ but can be improved by flipping a variable $x_{i}$. This flip can be realized by the permutation $\pi_{x_{i},r}$ where $r$ is $1$ if we want to flip $x_{i}$ from 0 to $1$ and $p_{i}-1$ if we want to flip it from $1$ to 0. Since flipping $x_{i}$ is an improving step, it additionally satisfies some clause $C$ and does not additionally violate any higher priority clause $C^{\prime}$; thus, applying $\pi_{x_{i},r}\circ\pi_{x_{j},s}$ moves the $1$ of $v$ away from the clause-violating position of $C$ but places no new $1$ on any clause-violating position of any higher-priority clause $C^{\prime}$. Also, it does not move a 1 onto any bad-color-position. Thus, $v$ is not a local minimum, either.
@@
We infer that every local minimum $v=s\circ\pi^{l}$ under the permutations $\pi^{e_{1}},\dots,\pi^{e_{M}}$ corresponds to a local maximum of the $3$-SAT/2-FLIP instance $F$. This concludes the proof of Theorem 5. ∎
@@ 8 Conclusion and Open Problems
We have shown that commutativity does not help when solving the orbit minimization problem, strengthening the result from [16].
@@
Our current reduction does not work for 3-SAT/1-FLIP because we need 3 variables per gate and a switching variable for the highest clauses. Hence, a new approach is needed here. As computing a global optimum is also NP-hard for 3-SAT with lexicographic weights (just reduce from the plain old decision problem of $3$-SAT), one might again suspect hardness. On the other hand, finding a globally optimal max cut on cubic graphs is NP-hard, while the local optima can be found in polynomial time[14, 11]. While $2$-SAT is tractable for lexicographic weights, it is harder than the unweighted case which is in NL, since our algorithm needs to remember a set of clauses that can be fulfilled so far. This is unlikely to improve:
@@
Another interesting PLS-problem with lexicographical weights is the travelling salesman problem. The usual approach does not work similarly as to $4$-SAT, since we would have to check whether a Hamiltonian cycle still exists without these edges. Recently, Heimann, Hoang and Hougardy investigated the complexity of this problem for general costs[6]. Their reduction however starts from Max-Cut, which is easy in the lexicographical setting. We believe that a reduction similar to the NP-completeness proof for Hamiltonian cycle would work, if the number of occurrences of each variable in the $k$-SAT instance is bounded by some constant.

### v2

@@ 1.7 Our results
The proof, given in Section 3, follows the approach of Krentel [DBLP:conf/focs/Krentel89, DBLP:conf/coco/Krentel89], who reduces from FLIP, but deviates from it in one crucial detail. With some additional work, we can show that $3$-SAT/2-FLIP with lexicographic payoffs is PLS-complete.
@@
We apply the previous results to answer an open question from Scheder and Tantow [SchederTantow2025]
@@
We prove the theorem in Section 6 The proof follows from a reduction very similar to Buchheim and Jünger [DBLP:journals/disopt/BuchheimJ05]. The permutations constructed in the reduction are commuting involutions (permutations of order 2) and thus the generated subgroup of the symmetric group is isomorphic to the additive group of $\mathbb{F}_{2}^{n}$, the vector space over $\mathbb{F}_{2}$.
We will give a second reduction, following the proof of Theorem 2 in [SchederTantow2025], where all permutations are powers of one single permutation $\pi$. In other words, we prove the following theorem:
@@
Finally, we extend our approach to game theory and the complexity of finding approximate Nash equilibria. This was proven to be PLS-complete in [DBLP:conf/stoc/SkopalikV08]. The hardness of lexicographic 4-SAT/1-Flip allows an easy proof and generalizes the original statement to exponential functions.
Theorem 6.
Finding $\alpha$-Nash equilibria in congestion games is PLS-complete for step functions and exponential functions as delay functions.
The proof is in Section 8
@@ 2 Krentel’s approach and why it does not directly work for lexicographic Max-$k$-SAT
Krentel proved this to be PLS-complete[DBLP:conf/coco/Krentel89] with the test circuit technique by reducing from FLIP. We sketch this proof again and show why it does not work for lexicographic weights. We are given a circuit $C$ with $n$ inputs and $m$ outputs. For the reduction, we imagine $n+1$ copies of $C$, called $C^{(0)},C^{(1)},\dots,C^{(n)}$. We want $C^{(0)}$ to compute $C(x)$ for the current input $x$, and each other $C^{(i)}$ ($1\leq i\leq n$) to compute $C(x\oplus\mathbf{e}_{i})$. For each circuit $C^{(i)}$ ($0\leq i\leq n$) we create a vector of $n$ input variables $\mathbf{x}^{(i)}=(x^{(i)}_{1},\dots,x^{(i)}_{n})$. We wish for $\mathbf{x}^{(i)}=\mathbf{x}^{(0)}\oplus\mathbf{e}_{i}$. For each circuit $i$ and gate $j$ we create a gate variable $g_{j}^{(i)}$. The variables involved in a gate can be consistent or inconsistent with the gate, and we call the gate correct or incorrect accordingly. We wish for all gates to be correct, but have to tolerate them being temporarily incorrect.
@@ 3 Lexicographic 4-SAT/2-Flip: Proof of Theorem 1
We have a switch variable $f$ (the needle in the above figure) that indicates whether TOP is considered the active circuit (if $f=1$) or BOTTOM is (if $f=0$). Finally, and this differs from [DBLP:conf/coco/Krentel89], we have a global output register $\vec{o}=(o_{1},o_{2},\dots,o_{m})$.
@@
A difficulty. Simulating a correct evaluation of all gates is relatively easy to do with $k$-SAT/1-FLIP. The problem arises when all circuits are correct, but the payoff output of the inactive circuit is better than that of the active output. We need a mechanism that allows us to make the inactive circuit active (and vice versa) while improving the payoff (the total value of satisfied clauses). In [DBLP:conf/coco/Krentel89] this is achieved by introducing clauses $(f\rightarrow v_{i}^{T})$ and $(\neg f\rightarrow v_{i}^{B})$, both of weight $2^{c-j}$ for some constant $c$. Making the BOTTOM active and TOP inactive (for example) will increase the total satisfied weight if and only if $\vec{v}^{B}>_{\rm lex}\vec{v}^{T}$. However, this only works if $(f\rightarrow v_{i}^{T})$ and $(\neg f\rightarrow v_{i}^{B})$ really have the same weight. If we insist on lexicographic weights and the first bit of both outputs is 0, say, and BOTTOM is active, then among $(f\rightarrow v_{1}^{T})$ and $(\neg f\rightarrow v_{1}^{B})$ only the first clause is satisfied, and it would never pay to make TOP active. This is where we need to introduce the global output and allow double flips.
@@ 6 Reduction to Abelian Polm—Proof of Theorem 4
We adapt a proof technique from [DBLP:journals/disopt/BuchheimJ05] to reduce $4$-SAT/1-FLIP to Abelian Polm. Let $F$ be our $4$-CNF formula, let $x_{1},\dots,x_{n}$ be its variables, and let $C_{1},\dots,C_{m}$ be its clauses.
@@ 7 Reduction to Cyclic Group Polm—Proof of Theorem 5
Conversely, if $v$ is a local minimum, then $v$ has no $1$ at any bad-color position and thus $v$ encodes, as described above, a truth assignment $\alpha$ to the variables $x_{1},\dots,x_{n}$. Suppose $\alpha$ is not locally optimal for $F$ but can be improved by flipping a variable $x_{i}$. This flip can be realized by the permutation $\pi_{x_{i},r}$ where $r$ is $1$ if we want to flip $x_{i}$ from 0 to $1$ and $p_{i}-1$ if we want to flip it from $1$ to 0. Since flipping $x_{i}$ is an improving step, it additionally satisfies some clause $C$ and does not additionally violate any higher priority clause $C^{\prime}$; thus, applying $\pi_{x_{i},r}$ moves the $1$ of $v$ away from the clause-violating position of $C$ but places no new $1$ on any clause-violating position of any higher-priority clause $C^{\prime}$. Also, it does not move a 1 onto any bad-color-position. Thus, $v$ is not a local minimum, either.
@@
We infer that every local minimum $v=s\circ\pi^{l}$ under the permutations $\pi^{e_{1}},\dots,\pi^{e_{M}}$ corresponds to a local maximum of the $4$-SAT/1-FLIP instance $F$. This concludes the proof of Theorem 5. ∎
8 An application to game theory
Lexicographic $k$-SAT can also help to show hardness for results, whose cost functions are in general not lexicographic. A congestion game can be represented as a tuple $(N,E,(d_{e})_{e\in E},(S_{i})_{i\in N})$, where $N$ is a set of players and $E$ is a set of resources. Each resource $e\in E$ has a delay $d_{e}:\mathbb{N}\to\mathbb{R}$ and each player $i$ has a set of strategies $S_{i}\subseteq 2^{E}$.
Each player $i$ plays a strategy $s_{i}$ that are combined in a strategy profile $s=(s_{1},s_{2},\dots,s_{n})$. The cost of such a strategy profile $s$ for a player $i$ is the sum of the delays of each resource in $s_{i}$, i.e. $C_{i}(s)=\sum_{e\in s_{i}}d_{e}(|\{j\in N\mid e\in s_{j}\}|)$. The players act selfishly and switch their strategy, if they can reduce their cost. A pure Nash-equilibrium is a strategy profile $s$ such that for any strategy profile $s^{\prime}$ that differs only in the strategy of player $i$, we have that $C_{i}(s)\leq C_{i}(s^{\prime})$, i.e. there is no reason to deviate alone. If a strategy is not a Nash equilibrium, then at least one player has an improving move that decreases his costs.
A simple reduction from local NAE-$3$-SAT to finding pure Nash equilibria is shown in [DBLP:conf/stoc/FabrikantPT04].
If the players are lazy and only change their strategy if their cost decreases a significant amount, then we ask for approximate Nash-equilibria. More formally, for a real $\alpha>1$, an $\alpha$-equilibrium is a strategy profile $s$ such that for any strategy profile $s^{\prime}$ that differs only in the strategy of player $i$, we have that $C_{i}(s)\leq\alpha\ C_{i}(s^{\prime})$. In [DBLP:conf/stoc/SkopalikV08], this was show to be PLS-complete for any computable $\alpha>1$ using an involved reduction from FLIP.
It is not possible to reuse the general reduction from $3$-SAT as there could be two clauses $C$ and $C^{\prime}$ with the same high weight $w$ and a variable $x$, such that $x\in C$ and $\neg x\in C^{\prime}$. Suppose now that in the 3-SAT instance, we flip the value of $x$ to improve a small weight $w^{\prime}\ll w$. Then this might not be an $\alpha$-improving move, as the change is only very little in comparison to the large weight $w$. These problems don’t occur with lexicographic weights.
Proposition 2 ([DBLP:conf/stoc/SkopalikV08]).
For any computable $\alpha>1$ and any congestion game $G$, finding an $\alpha$-equilibrium is PLS-complete.
Proof.
We reduce from lexicographic 4-SAT/1-FLIP. Let
$\phi=C_{1}\wedge C_{2}\wedge\dots\wedge C_{m}$
be a weighted 4-SAT formula with the clauses ordered from low-weight to high-weight (for example, clause $C_{j}$ having weight $2^{j}$ would do; as would any $c^{j}$ for $c\geq 2$); The set of resources are the clauses of $\phi$. We have for each variable $x$ in $\phi$ a player $x$ with the strategies $true=\{C\in\phi\mid\neg x\in C\}$ and $false=\{C\in\phi\mid x\in C\}$. Note that strategy profiles in the game are in a 1-to-1 correspondence with the Boolean assignments to the variables of $\phi$. The delay $d_{C}$ of clause/resource $C_{j}$ is defined to be zero if fewer than four players use it, and $(1+\alpha)^{j}$ if four players do (note that a resource can never be used by more than four players). Apart from the value $1+\alpha$, this is roughly as in [DBLP:conf/stoc/FabrikantPT04].
We must now show that a pure $\alpha$-Nash equilibrium corresponds to a local minimum of the 4-SAT/1-FLIP problem. Contrapositively, we show that a variable flip that improves the weight of satisfied clauses corresponds to a player that will change her strategy. Consider an improving flip of variable $x$ (from False to True, without loss of generality). Let $j$ be the highest position such that $C_{j}$ used to be unsatisfied before the flip and is now satisfied. The delay for player $x$ before the flip is therefore at least
$(1+\alpha)^{j}\ .$ (8.1)
Because our costs are lexicographic, the status of all higher clauses $C_{j+1},\dots,C_{m}$ does not change; in particular, all those that contain $\bar{x}$ are satisfied—before and after the flip. Thus, the delay for player $x$ after the flip is at most
$\sum_{i\leq j-1}(1+\alpha)^{i}=\frac{(1+\alpha)^{j}}{\alpha}\ .$ (8.2)
Thus, changing the strategy reduces the delay of player $x$ at least by a factor of $\alpha$ and player $x$ would change her strategy. ∎
In fact, we can even replace the step functions in the reduction by exponential functions: if $k$ players choose resource $C_{j}$, we define the delay to be
$(1+\alpha)^{km+j}\ .$ (8.3)
The term in (8.1) would now be $(1+\alpha)^{4m+j}$ and the one in (8.1) would be $\sum_{i\leq 4m+j-1}(1+\alpha)^{i}$ and thus the two terms would still be a factor of $\alpha$ apart. A technical issue arises when some of our clauses have fewer than four literals; in this case, we either have to multiply the cost function in (8.3) by $(1+\alpha)^{m(4-|C|)}$; alternatively, we could create up to three dummy players whose only strategy is to sit in short clauses and occupy those resources.
@@ 9 Conclusion and Open Problems
We have shown that commutativity does not help when solving the orbit minimization problem, strengthening the result from [SchederTantow2025].
@@
Our current reduction does not work for 3-SAT/1-FLIP because we need 3 variables per gate and a switching variable for the highest clauses. Hence, a new approach is needed here. As computing a global optimum is also NP-hard for 3-SAT with lexicographic weights (just reduce from the plain old decision problem of $3$-SAT), one might again suspect hardness. On the other hand, finding a globally optimal max cut on cubic graphs is NP-hard, while the local optima can be found in polynomial time[DBLP:journals/siamcomp/Poljak95, DBLP:conf/icalp/Loebl91]. While $2$-SAT is tractable for lexicographic weights, it is harder than the unweighted case which is in NL, since our algorithm needs to remember a set of clauses that can be fulfilled so far. This is unlikely to improve:
@@
Another interesting PLS-problem with lexicographical weights is the travelling salesman problem. The usual approach does not work similarly as to $4$-SAT, since we would have to check whether a Hamiltonian cycle still exists without these edges. Recently, Heimann, Hoang and Hougardy investigated the complexity of this problem for general costs[DBLP:conf/icalp/Heimann0H24]. Their reduction however starts from Max-Cut, which is easy in the lexicographical setting. We believe that a reduction similar to the NP-completeness proof for Hamiltonian cycle would work, if the number of occurrences of each variable in the $k$-SAT instance is bounded by some constant.
@@
Lexicographic k-SAT was useful to show hardness of finding approximate Nash equilibria for congestion games. An important subclass are (weighted or unweighted) congestion games where the delay at a resource is given by a polynomial in the number of players using this resource (resp. their total weight). We were not able to extend our results to this case. Note that in polynomial weighted congestion games finding $d^{\mathcal{O}(d)}$-equilibria is in PP[DBLP:journals/mor/GiannakopoulosN22, DBLP:journals/teco/CaragiannisFGS15], while equilibria are known to exists only for $\alpha\geq d$ where $d$ is the maximum degree of the polynomial functions[DBLP:journals/jcss/CaragiannisF21]. For $\alpha\in\mathcal{O}(\frac{d}{\log d})$ deciding if an equilibria exists is NP-hard[DBLP:journals/mor/ChristodoulouGGPW23].